2 identical balls of mass 'm' and carrying equal similar charges are suspended in air from the point by insulating threads of same length.when they are immersed in kerosene(dielectric constant=2,density=0.8gm/cc).the angle between the threads remains the same as before.find the density of the material of the blls?

Grade:12

1 Answers

Aman Bansal
12 years ago

Dear Kuldeep,

net in the y direction = 0

Tcos41 - mg = 0

Tcos41 = mg >>>>>>eq1



Fnet in the x direction = 0

Tsin41 - F = 0

Tsin41 = F >>>>>>eq 2



Now divide eq2 by eq1 to get:



tan41 = F / mg



F = (tan41)(mg)



And that's the Coulomb force. Then after you find the Coulomb force, apply Coulomb's law and solve for the charges:



F = kq^2 / r^2



F is the force solved up above. 

k is 9x10^9. 

r is the distance between the balls which can be obtained through a little trig **hint, the .4m will be the hypotenuse of 1 triangle post separation**



And then after plugging in the numbers, solve for q.



For the next part of the question, just relate tension to the equation used to solve for the Coulomb force above.



Tsin41 = Coulomb Force

T = F / sin41 

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert